# Ivona Najdenkoska    Mohammad Mahdi Derakhshani11footnotemark: 1    Yuki M. Asano Nanne van Noord    Marcel Worring    Cees G. M. Snoek22footnotemark: 2

### v1

@@ Abstract
We address the challenge of representing long captions in vision-language models, such as CLIP. By design these models are limited by fixed, absolute positional encodings, restricting inputs to a maximum of 77 tokens and hindering performance on tasks requiring longer descriptions. Although recent work has attempted to overcome this limit, their proposed approaches struggle to model token relationships over longer distances and simply extend to a fixed new token length. Instead, we propose a generalizable method, named TULIP, able to upgrade the token length to any length for CLIP-like models. We do so by improving the architecture with relative position encodings, followed by a training procedure that (i) distills the original CLIP text encoder into an encoder with relative position encodings and (ii) enhances the model for aligning longer captions with images. By effectively encoding captions longer than the default 77 tokens, our model outperforms baselines on cross-modal tasks such as retrieval and text-to-image generation.
@@ 1 Introduction
Obtaining text representations for long captions in vision-language models is an open research challenge. Within the context of text-only large language models, this problem has been studied extensively (Chung et al., 2024; Touvron et al., 2023; Bai et al., 2023; Gu & Dao, 2023; Su et al., 2024; Dubey et al., 2024; Pawar et al., 2024; Jiang et al., 2024; Lieber et al., 2024), yet, methods for context length expansion have only scarcely made their way into the vision-language domain. For instance, contrastive vision-language models like CLIP (Radford et al., 2021; Jia et al., 2021; Li et al., 2022) are constrained to short input sequences, capped at 77 tokens. Natively, such models based on Transformers (Vaswani et al., 2017) process their input as an unordered set of tokens and use positional encoding to preserve order information (Yun et al., 2019). In particular, CLIP models use absolute positional encodings, which rely on a predefined maximum number of tokens, thereby limiting the input sequence to 77 tokens by design. We introduce an efficient and generalizable method, TULIP, that upgrades the context window size of CLIP-like models.
@@
Our proposed method upgrades the context window size of CLIP-like models by utilizing relative positional encodings, which are not restricted to a fixed length. Changing the model in this manner normally requires expensive multi-modal retraining. Instead, we propose an adaptation phase to first distill the knowledge of the original CLIP into the new model with relative positional encodings using only caption data. Such a distillation approach is flexible and can be applied to CLIP-like models constrained by the standard 77-token window, transforming it into a model capable of handling longer captions. Afterwards, we perform full fine-tuning on the distilled model for a single epoch on 1 million image-caption pairs (Chen et al., 2023a), to further improve the alignment between images and long captions. After these phases, our TULIP model can ingest captions longer than the usual 77 tokens and we observe improved performance on tasks ranging from cross-modal retrieval to text-to-image generation, compared to (interpolated) fixed token-window baselines.
In addition to our new method, we introduce a new benchmark for long captions adapted from the recently introduced Dense Captioning Images (DCI) dataset (Urbanek et al., 2024). This benchmark overcomes the limitations of existing retrieval benchmarks which lack diversity, as they focus on specific scenes (e.g., Urban-1K (Zhang et al., 2024)), or are in-distribution datasets with already saturated performance (e.g., ShareGPT4v test set (Chen et al., 2023a)). Our results show that evaluating in a true long caption setting is crucial for evaluating contrastive vision-language mode as they unearth an increased performance gap as compared to prior benchmarks.
In summary, our contributions are: (1) We propose TULIP, the first contrastive vision-language model with relative positional encoding for long captions. (2) We propose a general training procedure to adapt the positional encoding of CLIP-like models using a two-step adaptation encompassing relative position distillation and expansion. (3) We demonstrate improved performance across different cross-modal retrieval and image generation tasks. We introduce a new benchmark Long-DCI for a more comprehensive evaluation of long-caption retrieval tasks.
@@ 2 Related Work
Position Encodings in Transformer Models. Transformers, the foundational architecture for many vision-language models, rely on positional encodings to compensate for the lack of inherent positional awareness in their set-based representation. In the natural language domain, absolute positional encodings (Vaswani et al., 2017) were initially proposed, where fixed embeddings are added to token embeddings based on their position in the sentence sequence. However, as models have grown larger and more complex, alternative approaches emerged such as relative positional encodings (Shaw et al., 2018; Press et al., 2021), randomized positional encodings (Ruoss et al., 2023), extrapolation techniques (Press et al., 2022) and positional interpolation (Chen et al., 2023b). Other works, such as Rotary Position Embedding (RoPE) (Su et al., 2024) and its variations (Chen et al., 2023b; Peng et al., 2023), apply relative positional encodings without any modifications to the self-attention mechanism, making it computationally efficient. Another recent approach is Contextual Position Encodings (CoPE) (Golovneva et al., 2024), which is a more general position encoding technique enabling one to attend to the $i$-th particular word, noun, or even sentence. In contrastive vision-language models, integrating positional information effectively across modalities remains an unsolved challenge and this is the primary focus of this paper.
@@
Contrastive Vision-Language Models with Long Captions. Contrastive Vision-language models have made significant strides in aligning visual and textual modalities, driven by the success of large-scale pre-trained models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), BLIP (Li et al., 2022) and many others (Garg et al., 2023; Vasu et al., 2024b; a; Cherti et al., 2023; Sun et al., 2023). These models leverage contrastive learning techniques to align image and text representations in a shared feature space, enabling robust performance on tasks such as image-text retrieval and zero-shot classification. However, all these models focus on short, global textual descriptions, often limited by small context windows, such as the 77-token limit in CLIP. Recent research has started considering this limitation, for instance, DCI (Urbanek et al., 2024) highlights the problem by pointing out that CLIP’s 77-token limit restricts the model from accommodating detailed, dense captions. They also introduce a human-annotated dataset containing dense image descriptions averaging above 1000 words each. Similarly, another work introduces DOCCI Onoe et al. (2024), a captioning dataset that provides detailed image descriptions that capture key challenges such as spatial relations, counting, text rendering, and world knowledge. DreamLIP (Zheng et al., 2024) studies the usage of synthetically generated long captions under the contrastive learning framework but uses subsampled short captions from longer captions instead of processing the complete long captions. Long-CLIP (Zhang et al., 2024) engages the challenge more directly, by stretching absolute position encodings through interpolation to enable fine-tuning on long captions. However, interpolation only partially addresses the limitations of absolute encodings when processing longer, complex captions. This is because interpolation merely extends the existing positional information without fundamentally altering its nature or capabilities. These limitations include diminished ability to capture fine-grained relative positions and poor generalization to longer captions (Pawar et al., 2024). We propose an alternative approach incorporating relative encodings without training from scratch, thereby enabling more effective processing of long captions and the comprehension of the pairwise token relationships therein.
@@ 3.3 Relative Position Distillation
This loss function ensures that the student model $s_{T}$ learns from the teacher model $f_{T}$, by leveraging the original model’s ability to align textual and visual information. After the distillation, the student model $s_{T}$ has the capabilities of the teacher model for encoding up to 77 tokens with relative positional encodings. The only missing feature is the ability to process longer context which is addressed next.
@@ 4 Experiments & Results
Datasets and downstream tasks. We evaluate TULIP on three downstream tasks: short caption cross-modal retrieval, long caption cross-modal retrieval, and text-to-image generation. For short caption cross-modal retrieval, we follow Zhang et al. (2024) and evaluate our model on the COCO2017 5k validation set (Lin et al., 2014) and the full Flickr30k dataset (Plummer et al., 2015). Similarly, for long caption cross-modal retrieval, we use two datasets namely ShareGPT4V test split and Urban-1K. Both datasets contain 1,000 image-caption pairs, collected by Zhang et al. (2024) and recaptioned using the ShareGPT4V captioning model.
It is important to recognize that long-caption cross-modal retrieval benchmarks come with several limitations. These include a lack of diversity, as some focus on narrowly defined scenes (e.g., Urban-1K), while others consist of in-distribution datasets where performance is already saturated (e.g., ShareGPT4V). To overcome these challenges, we introduce a new benchmark for long captions adapted from the recently introduced Dense Captioning Images (DCI) dataset (Urbanek et al., 2024). Long-DCI includes 7,000 human-annotated images and long caption pairs, with captions averaging 200 tokens per image. This human-led annotation process ensures diverse and accurate descriptions, avoiding the biases inherent in AI-generated captions like those from ShareGPT4V.
@@
Training details. Our training procedure comprises two phases: relative position distillation and relative position expansion, both utilize the ShareGPT4V dataset (Chen et al., 2023a) containing 1M image and long caption pairs. During the relative position distillation phase, we truncate captions to the first 77 tokens for both the teacher and student models. We train the student model using cosine loss as the distillation loss function for 20 epochs with a batch size of 640 using the AdamW optimizer (Loshchilov, 2017), setting the learning rate to 5e-4 with 1000 warmup steps. In the relative position expansion phase, we employ full-length captions without truncation, exposing the model to comprehensive textual details. The full TULIP model, featuring the new distilled text encoder, is fine-tuned using the NTK approach, with $\alpha$ empirically set to 8.0. For all main experiments, we use 248 number of tokens to match Long-CLIP’s context length for a fair comparison. Note that we can increase the length to more tokens, as shown in our ablations. We perform this fine-tuning stage for a single epoch with a batch size of 1280, a learning rate of 1e-5, and 1000 warmup steps using AdamW. We base our implementations on OpenAI’s pre-trained CLIP-ViT-B-16 and CLIP-ViT-L-14 architectures (Ilharco et al., 2021). Our code will be made publicly available.
@@
Table 2: Short caption cross-modal retrieval comparison on COCO and Flickr30k. TULIP shows competitive performance, often matching or exceeding Long-CLIP across different metrics and backbones.
@@ 4.1 Cross-modal Retrieval Comparison
We evaluate TULIP against the original CLIP, fine-tuned CLIP on ShareGPT4V (Chen et al., 2023a), and Long-CLIP Zhang et al. (2024) for both long-caption (Table 1) and short-caption (Table 2) cross-modal retrieval tasks. As shown in Table 1, our proposed model outperforms all benchmarks in long-caption cross-modal retrieval across 3 datasets on both image-to-text and text-to-image retrieval, utilizing two different vision backbones: ViT-B-16 and ViT-L-14 (Dosovitskiy et al., 2021). On the Long-DCI dataset, which is a more challenging benchmark for all approaches, we can observe that Long-CLIP already shows improvement over the original CLIP and that our TULIP model further improves on this. These results demonstrate the efficacy of TULIP in enhancing CLIP’s capabilities for long-caption cross-modal retrieval, particularly in diverse scenarios. For short captions, as shown in Table 2, we find that the tailored approach used by Long-CLIP for the first 20 tokens is beneficial for the short caption performance as they outperform CLIP even on short captions. Meanwhile, TULIP can obtain competitive performance without needing to tailor specifically to the first 20 tokens, which demonstrates the flexibility of relative positional encodings across different caption lengths. Overall, the results in Table1 and 2 indicate that TULIP is effective not only for long captions but also for maintaining competitive performance in short caption retrieval scenarios.
@@ 4.2 Text-to-Image Generation
In this section, we evaluate qualitatively how our proposed procedure enhances text-to-image generation by simply replacing the original CLIP ViT-L-14 text encoder with our TULIP model, in particular using the text encoder. We use Stable Diffusion XL (SDXL) (Podell et al., 2023) as the image generation backbone. Note that do not perform any additional training of the diffusion model. As observed in Figure 3, TULIP demonstrates improvements in both long and short caption understanding and modeling of nuanced details. For example, our version of SDXL + TULIP accurately depicts “the red tulip inside a wooden box” in the first example. This is not the case with SDXL-CLIP or SDXL-Long-CLIP which generate many tulips in the garden, missing the key detail of a single tulip being inside the box. This observation shows that our model can indeed capture finer details in the description. The second example shows that our model successfully generates “an old man sits on a rock” even when this detail appears beyond the 77-token limit. On the other hand, both base CLIP and Long-CLIP encoders fail at this. A similar observation is in the third example where the model correctly captures ‘people walking along the streets”. These examples are direct evidence that our approach can truly capture the meaning of words in longer captions. For shorter prompts, our model also shows enhanced caption understanding. For instance, it accurately represents the ferret with a “playful grin” demonstrating the enhanced association of attributes with their corresponding nouns. Even more impressively, it accurately places the ferret inside a “large glass jar”, showing correct spatial understanding - something both CLIP and Long-CLIP fail to achieve. These observations suggest that our TULIP text encoder enhances long caption comprehension, resulting in more accurate and contextually rich image generation across varying prompt lengths. We provide additional results in the appendix.
Figure 3: Text-to-Image Generation results. We replace the text encoder of SDXL with our own TULIP model. We observe improvements in both long captions understanding and capturing nuanced details, compared to the baselines CLIP-ViT-L-14 and Long-CLIP. Note that the // marks the 77-token boundary in the caption. The words in green indicate visual concepts that are correctly generated by TULIP and are missed by the baselines.

### v2

@@ Abstract
We address the challenge of representing long captions in vision-language models, such as CLIP. By design these models are limited by fixed, absolute positional encodings, restricting inputs to a maximum of 77 tokens and hindering performance on tasks requiring longer descriptions. Although recent work has attempted to overcome this limit, their proposed approaches struggle to model token relationships over longer distances and simply extend to a fixed new token length. Instead, we propose a generalizable method, named TULIP, able to upgrade the token length to any length for CLIP-like models. We do so by improving the architecture with relative position encodings, followed by a training procedure that (i) distills the original CLIP text encoder into an encoder with relative position encodings and (ii) enhances the model for aligning longer captions with images. By effectively encoding captions longer than the default 77 tokens, our model outperforms baselines on cross-modal tasks such as retrieval and text-to-image generation. The code repository is available at https://github.com/ivonajdenkoska/tulip.
@@ 1 Introduction
Obtaining text representations for long captions in vision-language models is an open research challenge. Within the context of text-only large language models, this problem has been studied extensively (Chung et al., 2024; Touvron et al., 2023; Bai et al., 2023; Gu & Dao, 2023; Su et al., 2024; Dubey et al., 2024; Pawar et al., 2024; Jiang et al., 2024; Lieber et al., 2024), yet, methods for context length expansion have only scarcely made their way into the vision-language domain. For instance, contrastive vision-language models like CLIP (Radford et al., 2021; Jia et al., 2021; Li et al., 2022) are constrained to short input sequences, capped at 77 tokens. Natively, such models based on Transformers (Vaswani et al., 2017) process their input as an unordered set of tokens and use positional encoding to preserve order information (Yun et al., 2019). In particular, CLIP models use absolute positional encodings, which rely on a predefined maximum number of tokens, thereby limiting the input sequence to 77 tokens by design. We introduce an efficient and generalizable method called TULIP, that upgrades the context window size of CLIP-like models.
@@
Our proposed method upgrades the context window size of CLIP-like models by utilizing relative positional encodings, which are not restricted to a fixed length. Changing the model in this manner normally requires expensive multi-modal retraining. Instead, we propose an adaptation phase to first distill the knowledge of the original CLIP into the new model with relative positional encodings using only caption data. Such a distillation approach is flexible and can be applied to CLIP-like models constrained by the standard 77-token window, transforming it into a model capable of handling longer captions. Afterwards, we perform full fine-tuning on the distilled model for a single epoch on 1 million image-caption pairs (Chen et al., 2023b), to further improve the alignment between images and long captions. After these phases, our TULIP model can ingest captions longer than the usual 77 tokens and we observe improved performance on tasks ranging from cross-modal retrieval to text-to-image generation, compared to (interpolated) fixed token-window baselines.
In addition to our new method, we introduce a new benchmark for long captions adapted from the recently introduced Dense Captioning Images (DCI) dataset (Urbanek et al., 2024). This benchmark overcomes the limitations of existing retrieval benchmarks which lack diversity, as they focus on specific scenes (e.g., Urban-1K (Zhang et al., 2024)), or are in-distribution datasets with already saturated performance (e.g., ShareGPT4v test set (Chen et al., 2023b)). Our results show that evaluating in a true long caption setting is crucial for evaluating contrastive vision-language mode as they unearth an increased performance gap as compared to prior benchmarks.
In summary, our contributions are: (1) We propose TULIP, the first contrastive vision-language model with relative positional encodings for long captions. (2) We propose a general training procedure to adapt the positional encodings of CLIP-like models using a two-step adaptation encompassing relative position distillation and expansion. (3) We demonstrate improved performance across different cross-modal retrieval and image generation tasks. We define a new benchmark Long-DCI for a more comprehensive evaluation of long-caption retrieval tasks.
@@ 2 Related Work
Position Encodings in Transformer Models. Transformers, the foundational architecture for many vision-language models, rely on positional encodings to compensate for the lack of inherent positional awareness in their set-based representation. In the natural language domain, absolute positional encodings (Vaswani et al., 2017) were initially proposed, where fixed embeddings are added to token embeddings based on their position in the sentence sequence. However, as models have grown larger and more complex, alternative approaches emerged such as relative positional encodings (Shaw et al., 2018; Press et al., 2021), randomized positional encodings (Ruoss et al., 2023), extrapolation techniques (Press et al., 2022) and positional interpolation (Chen et al., 2023c). Other works, such as Rotary Position Embedding (RoPE) (Su et al., 2024) and its variations (Chen et al., 2023c; Peng et al., 2023), apply relative positional encodings without any modifications to the self-attention mechanism, making it computationally efficient. Another recent approach is Contextual Position Encodings (CoPE) (Golovneva et al., 2024), which is a more general position encoding technique enabling one to attend to the $i$-th particular word, noun, or even sentence. In contrastive vision-language models, integrating positional information effectively across modalities remains an unsolved challenge and this is the primary focus of this paper.
@@
Contrastive Vision-Language Models with Long Captions. Contrastive Vision-language models have made significant strides in aligning visual and textual modalities, driven by the success of large-scale pre-trained models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), BLIP (Li et al., 2022) and many others (Garg et al., 2023; Vasu et al., 2024b; a; Cherti et al., 2023; Sun et al., 2023). These models leverage contrastive learning techniques to align image and text representations in a shared feature space, enabling robust performance on tasks such as image-text retrieval and zero-shot classification. However, all these models focus on short, global textual descriptions, often limited by small context windows, such as the 77-token limit in CLIP. Recent research has started considering this limitation, for instance, DCI (Urbanek et al., 2024) highlights the problem by pointing out that CLIP’s 77-token limit restricts the model from accommodating detailed, dense captions. Similarly, another work introduces DOCCI (Onoe et al., 2024), a captioning dataset that provides detailed image descriptions that capture key challenges such as spatial relations, counting, text rendering and world knowledge. DreamLIP (Zheng et al., 2024) studies the usage of synthetically generated long captions under the contrastive learning framework but uses subsampled short captions from longer captions instead of processing the complete long captions. Long-CLIP (Zhang et al., 2024) engages the challenge more directly, by stretching absolute position encodings through interpolation to enable fine-tuning with long captions. However, interpolation only partially addresses the limitations of absolute encodings when processing longer, complex captions. This is because interpolation merely extends the existing positional information without fundamentally altering its nature or capabilities. These limitations include diminished ability to capture fine-grained relative positions and poor generalization to longer captions (Pawar et al., 2024). We propose an alternative approach that incorporates relative encodings without training from scratch, thereby enabling more effective processing of long captions and the comprehension of the pairwise token relationships therein.
@@ 3.3 Relative Position Distillation
This loss function ensures that the student model $s_{T}$ learns from the teacher model $f_{T}$, by leveraging the original model’s ability to align image and text. After the distillation, the student model $s_{T}$ has the capabilities of the teacher model for encoding up to 77 tokens with relative positional encodings. The only missing feature is the ability to process longer context, which is addressed next.
@@ 4 Experiments & Results
Datasets and downstream tasks. We evaluate TULIP on three downstream tasks: short caption cross-modal retrieval, long caption cross-modal retrieval, and text-to-image generation. For short caption cross-modal retrieval, we follow Zhang et al. (2024) and evaluate our model on the COCO2017 5k validation set (Lin et al., 2014) and the full Flickr30k dataset (Plummer et al., 2015). Similarly, for long caption cross-modal retrieval, we use two datasets, namely ShareGPT4V test split and Urban-1K. Both datasets contain 1,000 image-caption pairs, collected by Zhang et al. (2024) and recaptioned using the ShareGPT4V captioning model.
It is important to recognize that long-caption cross-modal retrieval benchmarks come with several limitations. These include a lack of diversity, as some focus on narrowly defined scenes (e.g., Urban-1K), while others consist of in-distribution datasets where performance is already saturated (e.g., ShareGPT4V). To overcome these challenges, we define a new benchmark for long captions adapted from the recently introduced Dense Captioning Images (DCI) dataset (Urbanek et al., 2024). Long-DCI includes 7,000 human-annotated images and long caption pairs, with captions averaging 200 tokens per image. This human-led annotation process ensures diverse and accurate descriptions, avoiding the biases inherent in AI-generated captions like those from ShareGPT4V.
@@
Training details. Our training procedure comprises two phases: relative position distillation and relative position expansion, both utilize the ShareGPT4V dataset (Chen et al., 2023b) containing 1M image and long caption pairs. During the relative position distillation phase, we truncate captions to the first 77 tokens for both the teacher and student models. We train the student model using cosine loss as the distillation loss function for 20 epochs with a batch size of 640 using the AdamW optimizer (Loshchilov, 2017), setting the learning rate to 5e-4 with 1000 warmup steps. In the relative position expansion phase, we employ full-length captions without truncation, exposing the model to comprehensive-textual details. The full TULIP model, featuring the new distilled text encoder, is fine-tuned using the NTK approach, with $\alpha$ empirically set to 8.0. For all main experiments, we use 248 number of tokens to match Long-CLIP’s context length for a fair comparison. Note that we can increase the length to more tokens, as shown in our ablations. We perform this fine-tuning stage for a single epoch with a batch size of 1280, a learning rate of 1e-5, and 1000 warmup steps using AdamW. We base our implementations on OpenAI’s pre-trained CLIP-ViT-B-16 and CLIP-ViT-L-14 architectures (Ilharco et al., 2021).
@@
Table 2: Short caption cross-modal retrieval comparison on COCO and Flickr30k. TULIP shows competitive performance, often matching or exceeding Long-CLIP across different metrics and model backbones.
@@ 4.1 Cross-modal Retrieval Comparison
We evaluate TULIP against the original CLIP, fine-tuned CLIP on ShareGPT4V (Chen et al., 2023b), and Long-CLIP (Zhang et al., 2024) for both long-caption (Table 1) and short-caption (Table 2) cross-modal retrieval tasks. As shown in Table 1, our proposed model outperforms all benchmarks in long-caption cross-modal retrieval across 3 datasets on both image-to-text and text-to-image retrieval, utilizing two different vision backbones: ViT-B-16 and ViT-L-14 (Dosovitskiy et al., 2021).
On the Long-DCI dataset, which is a more challenging benchmark for all approaches, we can observe that Long-CLIP already shows improvement over the original CLIP and that our TULIP model further improves on this. These results demonstrate the efficacy of TULIP in enhancing CLIP’s capabilities for long-caption cross-modal retrieval, particularly in diverse scenarios. For short captions, as shown in Table 2, we find that the tailored approach used by Long-CLIP for the first 20 tokens is beneficial for the short caption performance as they outperform CLIP even on short captions. Whereas TULIP is able to obtain competitive performance without needing to specifically tailor to the first 20 tokens, which demonstrates the flexibility of relative positional encodings across different caption lengths. Overall, the results in Table 1 and 2 indicate that TULIP is effective not only for long captions but also for maintaining competitive performance in short caption retrieval scenarios.
@@ 4.2 Text-to-Image Generation
In this section, we evaluate qualitatively how our method enhances text-to-image generation by simply replacing the original CLIP ViT-L-14 text encoder in Stable Diffusion XL (SDXL) (Podell et al., 2023) with TULIP. Note that we do not perform any additional training of the diffusion model. As observed in Figure 3, TULIP demonstrates improvements in both long and short caption understanding and modeling of nuanced details, compared to T5-based models such as PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024), as well as CLIP-ViT-L-14 and Long-CLIP (Zhang et al., 2024). For example, our version of SDXL + TULIP accurately depicts “the red tulip inside a wooden box” in the first example. This is not the case with SDXL + CLIP or SDXL + Long-CLIP which generate many tulips in the garden, missing the key detail of a single tulip being inside the box. This observation shows that our model can indeed capture finer details in the description. The second example shows that our model successfully generates “an old man sits on a rock” even when this detail appears beyond the 77-token limit. On the other hand, base CLIP, Long-CLIP and PIXART-Alpha encoders fail at this. A similar observation is in the third example where the model correctly captures the “sharp suit” and “moustache” as very nuanced details, unlike the rest of the models. These examples are direct evidence that our approach can truly capture the meaning of words in longer captions. These observations suggest that our TULIP text encoder enhances long caption comprehension, resulting in more accurate and contextually rich image generation across varying prompt lengths. We provide additional image generation examples and human evaluation results in the appendix.
Figure 3: Text-to-Image Generation results. We replace the text encoder of SDXL with our own TULIP model. We observe improvements in capturing nuanced details, compared to T5-based models such as PIXART-Alpha and ELLA, as well as CLIP-ViT-L-14 and Long-CLIP. Note that the // marks the 77-token boundary in the caption. The words in green indicate visual concepts that are correctly generated by SDXL + TULIP and are missed by the baselines.